How many grams of 5% barium hydroxide solution will be needed to react 2 moles of sulfur (IV) oxide?


1
Expert's answer
2021-03-22T06:37:49-0400

Molar Mass of barium hydroxide = 171.34g/mol

5% of barium hydroxide = 8.567 g/mol

Sulfur IV oxide Molar Mass = 64.066g/mol

2 moles of sulfur IV oxide = 128.066g/mol

X×(1mol/128.066) × (2/1) = 2moles of sulfur

2X/128.066 = 2

2X = 128.066×2

X = 128.066

= 128.066 g of 5% barium hydroxide